OpenStudy (anonymous):

A new couch costs $2500. The sale advertised is 30% off. How much is the discount? (NOT THE FINAL PRICE) A. $600 B. $750 C. $1750

ganeshie8 (ganeshie8):

discount = 30 % of 2500 = 30/100 * 2500 = ?

OpenStudy (anonymous):

Divide the price by 100 to figure out what 1% is and multiply the result by 30.

OpenStudy (anonymous):

30% of $2500 is the discount. When you see "of" in a word problem, that means multiplication. So 30/100 * 2500= ?

OpenStudy (anonymous):

750

ganeshie8 (ganeshie8):

good work !

OpenStudy (anonymous):

thank you :) everyone!!

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

No worries.